OpenStudy (anonymous):

The functions f(x) and g(x) are shown below. f(x) = 0.05^x g(x) = (0.05)^x Which statement best describes the graph of f(x) and g(x)? The graph of f(x) will eventually exceed the graph of g(x). The graph of g(x) will eventually exceed the graph of f(x). The graphs will both have their y-intercept equal to 1. The graphs will both have their y-intercept equal to 0.05

OpenStudy (zzr0ck3r):

they are the same thing

OpenStudy (richyw):

\(a^0=1\)

OpenStudy (anonymous):

Oops wait lol. The first one isn't raised. It's just 0.05x the second is 0.05^x

OpenStudy (anonymous):

Okay I take that back lol

OpenStudy (richyw):

it was wrong anyways^

OpenStudy (anonymous):

Lol you know so whats the answer @richyw

OpenStudy (anonymous):

C

OpenStudy (richyw):

well think about what these graphs do as \(x\to\infty\). \[\lim_{x\to\infty}0.05x=\infty\]\[\lim_{x\to\infty}0.05^x=0\]

OpenStudy (anonymous):

Which line is f(x) and g(x)

OpenStudy (anonymous):

So A ?

OpenStudy (richyw):

f(x) is the linear one.

OpenStudy (anonymous):

Ohh okay. Lol you just taught me something and this wasn't even my question

OpenStudy (richyw):

yes I would say A.

OpenStudy (anonymous):

So A is right lol ?

OpenStudy (richyw):

well you haven't defined what A is... haha